# Energy Conditions and Quantum Information

### v1

@@ 1 Introduction
Any physically sensible system is expected to have positive definite energy, or at least a lower bound, to ensure a stable ground state. Also important in any physical theory is the law of the conservation of energy. In general relativity, the effects of matter fields on spacetime and conservation laws [^1] The energy-momentum tensor expresses a local conservation of energy and momentum but it does not immediately lead to global properties of energy and momentum of a matter field. In order to discuss the positivity and conservation laws of “total energy” of a system, one needs to introduce the notion of asymptotic flatness in order to define an isolated system (see Theorem 2.2.4 and Sec. 2.6 below). are captured by the “stress-energy-momentum tensor.” The positivity of energy density is then prescribed by the “energy conditions,” which are inequalities to be imposed on the energy-momentum tensor. As will be seen, depending on the choice of observers who measure energy density and also on the specific circumstances of their imposition, one can consider various different types of energy conditions, among which the most widely used are the weak (WEC), the dominant (DEC), the strong (SEC), and the null energy conditions (NEC) HE73 ; Wald84 . In other words, the choice of energy conditions is the selection of what one thinks physically reasonable matter fields are. In any case, roughly speaking, positive energy density generates the attractive nature of gravitation via the Einstein equations, leading to the focusing of families of world-lines of test particles or photons due to gravitational pull. This focusing effect—i.e., the convergence of causal geodesic curves—is key to understanding the physical consequences of gravitation and to establishing a number of important theorems in general relativity. For example, the singularity theorems of Hawking and Penrose Penrose:1964wq ; Hawking:1970zqf , which predict the existence of spacetime singularities inside a black hole or at the beginning of our universe, exploit certain energy conditions and geodesic focusing in an essential way in their proof. The laws of black hole thermodynamics are shown by using a combination of various energy conditions Hawking:1976de ; Wald:1999vt . Furthermore, by using the energy conditions, one can prohibit some exotic phenomena and structures from occurring, such as time-machines, wormholes, and faster-than-light travel Morris:1988tu ; in other words, such an exotic or pathological spacetime structure necessitates some sort of negative energy density.
@@
The positivity of energy of a physical material system is, however, not a trivial condition, especially when quantum effects are involved; there is no shortage of examples of negative energy in quantum mechanical systems. One of the most well-known examples is the Casimir effect. The Hawking radiation mentioned above also violates the positivity of energy. In fact, any locally (pointwisely) defined energy conditions can, in principle, be violated by quantum field effects [^2] It should also be noticed that even though local energy conditions can fail to hold under some quantum effects, they often admit lower bounds, hence satisfying what is called “quantum energy inequality.” See, e.g., Yurtsever:1994wc ; Fewster:2012yh ; Kontou:2020bta for quantum energy inequalities and their relations to the (averaged) energy conditions. . For this reason, various types of non-local modifications of energy conditions have been proposed, such as those by taking the average of a local energy condition along an observer’s world-line (called the averaged energy conditions), or by including non-local quantum field theoretic quantities, e.g., entanglement entropy of quantum fields (called the quantum null energy conditions). Such non-locally modified energy conditions turn out to play a key role in connecting quantum information with spacetime geometry.
@@
The purpose of this article is to review the basics of energy conditions in general relativity, their uses in gravitational physics, their recent applications in quantum field theory and holography, and their relations to quantum information. This article is not intended to cover the entire range of relevant subjects, to be comprehensive, or to provide a complete list of existing literature. Instead, we will introduce key ideas and methods of the uses of energy conditions and quantum information, by focusing on some specific topics. In the first part, we mainly focus on the four locally defined energy conditions (WEC, DEC, SEC, and NEC), with a particular emphasis on their role in several theorems in general relativity. In the second part, we first discuss how quantum effects lead to the violation of these locally defined energy conditions. Then we review some non-local modifications of the above mentioned energy conditions (ANEC and QNEC). We also discuss the role of (A)NEC in quantum field theory and in the context of the AdS/CFT correspondence.
@@
This article is organized as follows. In the next section, we discuss various types of the pointwise energy conditions and their uses in the context of classical theory of general relativity. In particular, we describe how the energy conditions are used in the singularity theorems, black hole mechanics and topology theorems. Most of the discussion in Sec. 2 is based on Hawking-Ellis HE73 and Wald Wald84 . We also discuss the violation of the pointwise energy conditions and introduce the averaged energy conditions, such as ANEC. A number of examples of exotic spacetime structures, such as wormholes, superluminal travel, and their connection to the violation of the energy conditions can be found in literature, e.g., Lobo:2017cay . More complete list of existing literature on the violation of locally defined, as well as averaged energy conditions can be found in, e.g., Fewster:2012yh ; Kontou:2020bta ; Martin-Moruno:2017exc . In Sec. 3, we review the energy conditions and their derivations in quantum field theory in Minkowski spacetime. We begin with the basic notions of quantum information such as the reduced density matrix and modular Hamiltonian for a subsystem, and introduce entanglement entropy and relative entropy as measures of quantum information. We derive the Bekenstein bound and ANEC from the positivity and monotonicity of relative entropy by following Casini:2008cr and Faulkner:2016mzt , respectively. We also outline the derivation of quantum null energy conditions (QNEC) by using relative entropy in Ceyhan:2018zfg . In Sec. 4, we review the connection between the energy conditions and the holographic principle. In particular, we explain the black hole information paradox in more detail in the context of the AdS/CFT correspondence, introducing the recent advancements such as the holographic entanglement entropy or the Ryu-Takayanagi formula Ryu:2006bv ; Ryu:2006ef , quantum extremal surface (QES) Faulkner:2013ana ; Engelhardt:2014gca , and Island formula Almheiri:2019hni ; Penington:2019npb ; Almheiri:2019psf . Then, we discuss the holographic proof of ANEC and its generalization by focusing on a role of causality in the AdS/CFT correspondence. Section 5 is devoted to summary and discussions.
@@ 2 Energy conditions in general relativity
In this section, we will review basic properties of the energy conditions and their applications in general relativity and gravitational physics, including the singularity theorems, the laws of black hole mechanics, and the topological censorship. Our main focus is on the uses of four main, locally (pointwisely) defined energy conditions, but we also discuss their violations and non-local generalizations at the end of this section.
@@ 2.1 The Einstein equations
Spacetime geometry can be described in terms of the pair of the metric tensor $g_{\mu\nu}$ with Lorentzian signature and a manifold $M$. The Einstein equations are a set of second-order differential equations for the metric tensor, expressed as
@@
where $R_{\mu\nu}$ and $R$ are the Ricci tensor and scalar curvature, respectively, constructed from the Riemann tensor333 The Riemann tensor is defined as follows: $R^{\lambda}{}_{\sigma\mu\nu}=\partial_{\mu}\Gamma^{\lambda}_{\sigma\nu}-\partial_{\nu}\Gamma^{\lambda}_{\sigma\mu}+\Gamma^{\lambda}_{\alpha\mu}\Gamma^{\alpha}_{\sigma\nu}-\Gamma^{\lambda}_{\alpha\nu}\Gamma^{\alpha}_{\sigma\mu}\,$, (2) where $\Gamma^{\lambda}_{\mu\nu}=\dfrac{1}{2}g^{\lambda\sigma}(\partial_{\mu}g_{\sigma\nu}+\partial_{\nu}g_{\sigma\mu}-\partial_{\sigma}g_{\mu\nu})$ is the Christoffel symbol and $R_{\mu\nu}:=R^{\alpha}{}_{\mu\alpha\nu}$ and $R:=g^{\mu\nu}R_{\mu\nu}$. For geometric quantities, we follow the convention and notations of Wald’s book Wald84 . . On the right-hand side are the energy-momentum (or stress-energy) tensor $T_{\mu\nu}$ for cosmic matter fields and $G$ denotes Newton’s gravitational constant. Due to the Bianchi-identity the divergence of the left-hand side is automatically vanishing, and accordingly, the divergence of the right-hand side is also required to satisfy,
@@
More generally, when an action $S_{M}$ for a matter field is given, the energy-momentum tensor $T_{\mu\nu}$ is obtained by taking a variation of $S_{M}$ with respect to the metric tensor:
@@
$T^{(\phi)}_{\mu\nu}=\dfrac{-2}{\sqrt{-g}}\dfrac{\delta S_{\phi}}{\delta g^{\mu\nu}}=\nabla_{\mu}\phi\nabla_{\nu}\phi-\dfrac{1}{2}g_{\mu\nu}\left[g^{\alpha\beta}\nabla_{\alpha}\phi\nabla_{\beta}\phi+2V(\phi)\right]\,.$ (7)
@@ 2.2 Pointwise energy conditions
The positivity of energy can, in principle, depend on an observer who measures energy and momentum. Besides the positivity, also important is the causality; in relativistic theory, energy and momentum associated with a particle form a single four-vector (four-momentum $p^{\mu}$), which is required to be timelike or null. Depending upon the choice of an observer, one may consider different types of energy conditions on the given energy-momentum tensor, as will be seen below. However, before going into more details of possible energy conditions, we should discuss why we need to impose them.
@@
    The first reason is that if no restrictions are imposed on the energy-momentum tensor (other than it satisfying conservation laws), virtually any metric could be considered a solution to the Einstein equations. This is because for any given metric, one can simply view the Einstein equations as the defining equations for an energy-momentum tensor. However, such an arbitrary definition would not represent a physically meaningful spacetime. Instead, we must first impose realistic conditions on the matter fields. Only then should the Einstein equations be used to determine a physically valid metric.
@@
{itembox}
[l]Definition 2.2.1 (The Weak Energy Condition (WEC))
$T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$ for arbitrary timelike vector $\xi^{\mu}$.
@@
{itembox}
[l]Definition 2.2.2 (The Dominant Energy Condition (DEC)) For arbitrary future-directed timelike vector $\xi^{\mu}$, $T^{\mu}{}_{\nu}\xi^{\nu}$ becomes a future-directed timelike or null vector.
@@
    {itembox}
    [l]Proposition 2.2.3 (Conservation law) Suppose the energy-momentum tensor for matter fields $T_{\mu\nu}$ satisfies the local conservation $\nabla_{\nu}T^{\mu\nu}=0$ and the DEC. If $T_{\mu\nu}=0$ on a closed, achronal hypersurface $\Sigma$, then $T_{\mu\nu}=0$ everywhere in the domain of dependence $D(\Sigma)$.
@@
  • Another important consequence of the DEC is the positivity of the total energy of an isolated system. In general relativity, an isolated system is defined in terms of asymptotic flatness, and the total energy defined at spatial infinity is called the “Arnowitt-Deser-Misner (ADM)” energy Arnowitt:1962hi , whose positivity was shown by Schoen-Yau Schon:1981vd , Witten Witten:1981mf : {itembox}[l]Theorem 2.2.4 (Positive energy theorem) The total energy is positive for initial data on an asymptotically flat Cauchy surface, provided the DEC holds.
@@
{itembox}
[l]Definition 2.2.5 (The Strong Energy Condition (SEC))
$\left(T_{\mu\nu}-\frac{1}{2}T^{\lambda}{}_{\lambda}g_{\mu\nu}\right)\xi^{\mu}\xi^{\nu}\geqslant 0\quad\mbox{for arbitrary timelike vector $\xi^{\mu}$.}$
@@
  • There is no containment relationship between SEC and WEC. For example, for a negative cosmological constant, $\Lambda<0$, SEC is satisfied, but WEC fails to hold.
{itembox}
[l]Definition 2.2.6 (The Null Energy Condition (NEC))
$T_{\mu\nu}k^{\mu}k^{\nu}\geqslant 0$ for arbitrary null vector $k^{\mu}$.
@@
  • NEC is equivalent to WEC with a timelike vector $\xi^{\mu}$ replaced by a null vector $k^{\mu}$. If WEC holds, then by continuity, NEC also holds. Similarly, if SEC conditions holds, then for any $\xi^{\mu}$, $T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant(1/2)T^{\lambda}{}_{\lambda}\xi_{\mu}\xi^{\nu}$, and hence, by continuity, NEC also holds.
  • The violation of NEC (NCC) implies the violation of WEC. Suppose NEC is violated. Then, for any future-directed null vector $k^{\mu}$, there exists a constant $\delta>0$ such that
@@
    Let $u^{\mu}$ be an appropriate future-directed timelike vector. Then, $\xi^{\mu}:=k^{\mu}+\epsilon u^{\mu},\,(\epsilon>0)$ can also be a timelike vector. For this timelike vector $\xi^{\mu}$,
@@
Table 1: Four locally defined energy conditions. Here $\xi^{\mu}$ and $k^{\mu}$ denote any timelike and null vector, respectively, and “causality restriction” means that superluminal flow of energy is not allowed.
Definition Purpose and effects Type I case
WEC    $T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$ positivity restriction    $\rho\geqslant 0$, $\rho+p_{i}\geqslant 0$
DEC $T_{\mu\nu}\xi_{1}^{\mu}\xi_{2}^{\nu}\geqslant 0$    positivity and causality restriction    $\rho\geqslant|p_{i}|\geqslant 0$

### v2

@@ 1 Introduction
Any physically sensible system is expected to have positive definite energy, or at least a lower bound, to ensure a stable ground state. Also important in any physical theory is the law of the conservation of energy. In general relativity, the effects of matter fields on spacetime and conservation laws [^1] The energy-momentum tensor allows us to express a local conservation of energy and momentum but it does not immediately lead to global properties of energy and momentum of a matter field. In order to discuss the positivity and conservation laws of “total energy” of a system, one needs to introduce the notion of asymptotic flatness in order to define an isolated system (see Theorem 2.2.4 and Sec. 2.6 below). are captured by the “stress-energy-momentum tensor.” The positivity of energy density is then prescribed by the “energy conditions,” which are inequalities to be imposed on the energy-momentum tensor. As will be seen, depending on the choice of observers who measure energy density and also on the specific circumstances of their imposition, one can consider various different types of energy conditions, among which the most widely used are the weak (WEC), the dominant (DEC), the strong (SEC), and the null energy conditions (NEC) HE73 ; Wald84 . In other words, the choice of energy conditions is the selection of what one thinks physically reasonable matter fields are. In any case, roughly speaking, positive energy density generates the attractive nature of gravitation via the Einstein equations, leading to the focusing of families of world-lines of test particles or photons due to gravitational pull. This focusing effect—i.e., the convergence of causal geodesic curves—is key to understanding the physical consequences of gravitation and to establishing a number of important theorems in general relativity. For example, the singularity theorems of Hawking and Penrose Penrose:1964wq ; Hawking:1970zqf , which predict the existence of spacetime singularities inside a black hole or at the beginning of our universe, exploit certain energy conditions and geodesic focusing in an essential way in their proof. The laws of black hole thermodynamics are shown by using a combination of various energy conditions Hawking:1976de ; Wald:1999vt . Furthermore, by using the energy conditions, one can prohibit some exotic phenomena and structures from occurring, such as time-machines, wormholes, and faster-than-light travel Morris:1988tu ; in other words, such an exotic or pathological spacetime structure necessitates some sort of negative energy density.
@@
The positivity of energy of a physical material system is, however, not a trivial condition, especially when quantum effects are involved; there is no shortage of examples of negative energy in quantum mechanical systems. One of the most well-known examples is the Casimir effect, which is experimentally well-established (see, e.g., Klimchitskaya:2009cw and references therein). The Hawking radiation mentioned above also violates the positivity of energy. In fact, any locally (pointwisely) defined energy condition can, in principle, be violated by quantum field effects [^2] It should also be noticed that even though local energy conditions can fail to hold under some quantum effects, they often admit lower bounds, hence satisfying what is called “quantum energy inequality.” See, e.g., Yurtsever:1994wc ; Fewster:2012yh ; Kontou:2020bta for quantum energy inequalities and their relations to the (averaged) energy conditions. . For this reason, various types of non-local modifications of energy conditions have been proposed, such as those by taking the average of a local energy condition along an observer’s world-line (called the averaged energy conditions), or by including non-local quantum field theoretic quantities, e.g., entanglement entropy of quantum fields (called the quantum null energy conditions). Such non-locally modified energy conditions turn out to play a key role in connecting quantum information with spacetime geometry.
@@
The purpose of this article is to review the basics of energy conditions in general relativity, their uses in gravitational physics, their recent applications in quantum field theory, and relations to quantum information. This article is not intended to cover the entire range of relevant subjects, to be comprehensive, or provide a complete list of existing literature. Instead, we will introduce key ideas and methods of the uses of energy conditions and quantum information, by focusing on some specific topics. In the first part, we mainly focus on the four locally defined energy conditions (WEC, DEC, SEC, NEC), with a particular emphasis on their role in several theorems in general relativity. In the second part, we first discuss how quantum effects lead to the violation of these locally defined energy conditions. Then we review some non-local modifications of the above mentioned energy conditions (ANEC and QNEC). We also discuss the role of (A)NEC in quantum field theory and in the context of the AdS/CFT correspondence.
@@
This article is organized as follows. In the next section, we discuss various types of the pointwise energy conditions and their uses in the context of classical theory of general relativity. In particular, we describe how the energy conditions are used in the singularity theorems, black hole mechanics and topology theorems. Most of the discussion in Sec. 2 is based on Hawking-Ellis HE73 and Wald Wald84 . We also discuss the violation of the pointwise energy conditions and introduce the averaged energy conditions, such as ANEC. A number of examples of exotic spacetime structures, such as wormholes, superluminal travel, and their connection to the violation of the energy conditions can be found in literature, e.g., Lobo:2017cay . More complete list of existing literature on the violation of locally defined, as well as averaged energy conditions can be found in, e.g., Fewster:2012yh ; Kontou:2020bta ; Martin-Moruno:2017exc . In Sec. 3, we review the energy conditions and their derivations in quantum field theory in Minkowski spacetime. We begin with the basic notions of quantum information such as the reduced density matrix and modular Hamiltonian for a subsystem, and introduce entanglement entropy and relative entropy as measures of quantum information. We derive the Bekenstein bound and ANEC from the positivity and monotonicity of relative entropy by following Casini:2008cr and Faulkner:2016mzt , respectively. We also outline the derivation of quantum null energy conditions (QNEC) by using relative entropy in Ceyhan:2018zfg . In Sec. 4, we review the connection between the energy conditions and the holographic principle. In particular, we explain the black hole information paradox in more detail in the context of the AdS/CFT correspondence, introducing the recent advancements such as the holographic entanglement entropy or the Ryu-Takayanagi formula Ryu:2006bv ; Ryu:2006ef , quantum extremal surface (QES) Faulkner:2013ana ; Engelhardt:2014gca , and Island formula Almheiri:2019hni ; Penington:2019npb ; Almheiri:2019psf . Then, we discuss the holographic proof of ANEC and its generalization by focusing on the role of causality in the AdS/CFT correspondence. Section 5 is devoted to summary and discussions.
@@ 2 Energy conditions in general relativity
In this section, we will review basic properties of the energy conditions and their applications in general relativity and gravitational physics, including the singularity theorems, the laws of black hole mechanics, and the topological censorship. Our main focus is on the uses of four main, locally (pointwisely) defined energy conditions, but we also discuss their violations and non-local generalizations at the end of this section. Unless specifically stated, the assertions, theorems, and formulas discussed in this section assume a four-dimensional spacetime. Most of these formulas and theorems remain valid in general dimensions, either without modification or with obvious dimensional extensions.
@@ 2.1 The Einstein equations
Spacetime geometry can be described in terms of the pair of a metric tensor $g_{\mu\nu}$ with Lorentzian signature and a connected manifold $M$. The Einstein equations are a set of second-order differential equations for the metric tensor, expressed as
@@
where $R_{\mu\nu}$ and $R$ are the Ricci tensor and scalar curvature, respectively, constructed from the Riemann tensor [^3] The Riemann tensor is defined as follows: $R^{\lambda}{}_{\sigma\mu\nu}=\partial_{\mu}\Gamma^{\lambda}_{\sigma\nu}-\partial_{\nu}\Gamma^{\lambda}_{\sigma\mu}+\Gamma^{\lambda}_{\alpha\mu}\Gamma^{\alpha}_{\sigma\nu}-\Gamma^{\lambda}_{\alpha\nu}\Gamma^{\alpha}_{\sigma\mu}\,$, (2) where $\Gamma^{\lambda}_{\mu\nu}=\dfrac{1}{2}g^{\lambda\sigma}(\partial_{\mu}g_{\sigma\nu}+\partial_{\nu}g_{\sigma\mu}-\partial_{\sigma}g_{\mu\nu})$ is the Christoffel symbol and $R_{\mu\nu}:=R^{\alpha}{}_{\mu\alpha\nu}$ and $R:=g^{\mu\nu}R_{\mu\nu}$. For geometric quantities, we follow the convention and notations of Wald’s book Wald84 . . On the right-hand side are the energy-momentum (or stress-energy) tensor $T_{\mu\nu}$ for cosmic matter fields and $G$ denotes Newton’s gravitational constant. Due to the Bianchi-identity the divergence of the left-hand side is automatically vanishing, and accordingly, the divergence of the right-hand side is also required to satisfy,
@@
In general, when an action $S_{M}$ for a matter field is given, the energy-momentum tensor $T_{\mu\nu}$ is obtained by taking a variation of $S_{M}$ with respect to the metric tensor:
@@
$T^{(\phi)}_{\mu\nu}$ $=$ $\dfrac{-2}{\sqrt{-g}}\dfrac{\delta S_{\phi}}{\delta g^{\mu\nu}}$ (7)
$=$ $\nabla_{\mu}\phi\nabla_{\nu}\phi-\dfrac{1}{2}g_{\mu\nu}\left[g^{\alpha\beta}\nabla_{\alpha}\phi\nabla_{\beta}\phi+2V(\phi)\right]\,.$
@@ 2.2 Pointwise energy conditions
The positivity of energy can, in principle, depend on an observer who measures energy and momentum. Besides the positivity, also important is the causality; in relativistic theory, energy and momentum associated with a particle form a single four-vector (four-momentum $p^{\mu}$), which is required to be timelike or null, and also be future pointing. Depending upon the choice of an observer, one may consider different types of energy conditions on the given energy-momentum tensor, as will be seen below. However, before going into more details of possible energy conditions, we should discuss why we need to impose them.
@@
    The first reason is that if no restrictions are imposed on the energy-momentum tensor (other than it satisfying conservation laws), virtually any metric could be considered a solution to the Einstein equations. This is because for any given metric, one can simply view the Einstein equations as the defining equations for an energy-momentum tensor [^5] This procedure for defining an energy-momentum tensor is sometimes known as the “Synge G-method” or the “Nariai-method” (see, e.g., Ellis:2023css ). . However, such an arbitrary definition would not represent a physically meaningful spacetime. Instead, we must first impose realistic conditions on the matter fields. Only then should the Einstein equations be used to determine a physically valid metric.
@@
  Definition 2.2.1 (Weak Energy Condition)   $T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$ for arbitrary timelike vector $\xi^{\mu}$.
@@
  Definition 2.2.2 (Dominant Energy Condition)   For arbitrary future-directed timelike vector $\xi^{\mu}$, $T^{\mu}{}_{\nu}\xi^{\nu}$ becomes a future-directed timelike or null vector.
@@
      Proposition 2.2.3 (Conservation law)   Suppose the energy-momentum tensor for matter fields $T_{\mu\nu}$ satisfies the local conservation $\nabla_{\nu}T^{\mu\nu}=0$ and the DEC. If $T_{\mu\nu}=0$ on a closed, achronal hypersurface $\Sigma$, then $T_{\mu\nu}=0$ everywhere in the domain of dependence $D(\Sigma)$.
@@
  • Another important consequence of the DEC is the positivity of the total energy of an isolated system. In general relativity, an isolated system is defined in terms of asymptotic flatness, and the total energy defined at spatial infinity is called the “Arnowitt-Deser-Misner (ADM)” energy Arnowitt:1962hi , whose positivity was shown by Schoen-Yau Schon:1981vd , Witten Witten:1981mf :
      Theorem 2.2.4 (Positive energy theorem)   The total energy is positive for initial data on an asymptotically flat Cauchy surface, provided the DEC holds.
@@
  Definition 2.2.5 (Strong Energy Condition)   $\left(T_{\mu\nu}-\dfrac{1}{2}T^{\lambda}{}_{\lambda}g_{\mu\nu}\right)\xi^{\mu}\xi^{\nu}\geqslant 0$ for arbitrary timelike vector $\xi^{\mu}$.
@@
  • SEC does not imply WEC. For example, for a negative cosmological constant, $\Lambda<0$, SEC is satisfied, but WEC fails to hold.
  Definition 2.2.6 (Null Energy Condition)   $T_{\mu\nu}k^{\mu}k^{\nu}\geqslant 0$ for arbitrary null vector $k^{\mu}$.
@@
  • NEC is analogous to WEC, with a timelike vector $\xi^{\mu}$ replaced by a null vector $k^{\mu}$. If WEC holds, then by continuity, NEC also holds. Similarly, if SEC conditions holds, then for any $\xi^{\mu}$, $T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant(1/2)T^{\lambda}{}_{\lambda}\xi_{\mu}\xi^{\nu}$, and hence, by continuity, NEC also holds.
  • The violation of NEC (NCC) implies the violation of WEC. Suppose NEC is violated. Then, for some future-directed null vector $k^{\mu}$, there exists a constant $\delta>0$ such that
@@
    Let $u^{\mu}$ be an appropriate future-directed timelike vector. Then, $\xi^{\mu}:=k^{\mu}+\epsilon u^{\mu},\,(\epsilon>0)$ is also a timelike vector. For this timelike vector $\xi^{\mu}$,
@@
Table 1: Four locally defined energy conditions. Here $\xi^{\mu}$ and $k^{\mu}$ denote any timelike and null vector, respectively, and “causality” means that superluminal flow of energy is not allowed. The timelike vectors $\xi_{1}^{\mu},\>\xi_{2}^{\nu}$ appearing in DEC are assumed to share the same time orientation.
Definition Effects Type I case
WEC    $T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$ positivity    $\rho\geqslant 0$, $\rho+p_{i}\geqslant 0$
DEC $T_{\mu\nu}\xi_{1}^{\mu}\xi_{2}^{\nu}\geqslant 0$    positivity and causality    $\rho\geqslant|p_{i}|\geqslant 0$